# Statistical Taylor Expansion111

### v4

@@ 2.3 Statistical Taylor Expansion
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines bounded momentum $\zeta(n)$, in which $0<\varrho,\sigma$ are bounding factors. Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@ 2.4 Statistical Bounding
When the bounding factors $\varrho$ and $\sigma$ are limited, Formula (2.4) provides statistical bounding $\tilde{x}\in[x-\varrho\delta x,x+\sigma\delta x]$. The probability $\epsilon$ for $\tilde{x}\not\in[x-\varrho\delta x,x+\sigma\delta x]$ is defined as bounding leakage.
@@ 2.5 Bounding Symmetry and Asymptote
A bounding $[x-\varrho\delta x,x+\sigma\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. To achieve mean-preserving, $\sigma$ determines $\varrho$. With mean-preserving bounding, Formula (2.11) and (2.12) give the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
@@
For any input distribution $\rho(\tilde{x},x,\delta x)$ symmetric around its mean $x$, any $[x-\sigma\delta x,x+\sigma\delta x]$ bounding has property $\zeta(2n+1)=0$, which simplifies statistical Taylor expansion further.
Empirically, as Formula (2.15), (2.18), and (2.23) will demonstrate, when $n\rightarrow+\infty$, $\zeta(n)\rightarrow\sigma^{n}/n$.
@@ 2.5.1 Uniform Input Uncertainty
For uniform distribution, the bounding factor is $\sigma=\sqrt{3}$, and the bounding leakage $\epsilon=0$. According to Formula (2.15), $\zeta(2n)=(\sqrt{3})^{2n}/(2n+1)$ while $\zeta(2n+1)=0$.
@@ 2.5.2 Gaussian Input Uncertainty
If $\sigma=\infty$, $\zeta(2n)=(2n-1)!!$!, which causes Formula (2.30) for $\log(x\pm\delta x)$ and Formula (2.33) for $(x\pm\delta x)^{c}$ to diverge. When $\sigma$ has an upper bound, Formula (2.4) becomes Formula (2.16):
$\zeta(2n)=$ $2N(\sigma)\sigma^{2n}\sum_{j=1}^{\infty}\sigma^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!};$ (2.16)
$=$ $(2n-1)\zeta(\sigma,2n-2)-2N(\sigma)\sigma^{2n-1};$ (2.17)
$\sigma^{2}\ll 2n:\;\;\;\zeta(2n)\simeq 2N(\sigma)\frac{\sigma^{2n+1}}{2n+1};$ (2.18)
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$! according to Formula (2.17). When $\sigma=5,n<5$, $\left|1-\eta(2n)/(2n-1)!!\right|<10^{-3}$.
  • For large $2n$, Formula (2.16) reduces to Formula (2.18), which shows that $\zeta(2n)$ increases slower than $\sigma^{2n}$ for increasing $2n$.
Formula (2.19) gives the bounding leakage $\epsilon$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5].
$\epsilon=2-2\xi(\frac{1}{2}+\sigma);$ (2.19)
@@ 2.5.3 An Input Uncertainty with Limited Range
Formula (2.20) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\varrho\lambda,\sigma\lambda]$ satisfying $0<\varrho<1<\sigma$. Formula (2.21) shows its mean-preserving equation. Formula (2.22) shows its bounded momentum. Formula (2.23) shows its asymptote. Formula (2.24) shows its bounding leakage.
@@
$e^{-\varrho}\varrho^{2}=e^{-\sigma}\sigma^{2};$ (2.21)
$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\sigma}\sum_{j=0}^{n+1}\frac{\sigma^{j}}{j!}\right);$ (2.22)
$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\sigma^{n+2}}{n+2}e^{-\sigma};$ (2.23)
$\epsilon$ $=1-e^{-\varrho}+e^{-\sigma};$ (2.24)
@@ 3.1 Choices of Bounding Factors
For discussion simplicity, only Gaussian input uncertainty will be considered in this paper. For Gaussian input uncertainty, when $\sigma=5$, $\epsilon=5.73\times 10^{-7}$ according to Formula (2.19), which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3]. Thus, $\sigma=5$ in variance arithmetic for Gaussian input uncertainty.
@@ 3.2 Low-Order Approximation
When $\sigma=5,n<5$, Formula (2.27), and (2.30) can be simplified as Formula (3.1), and (3.2), respectively.
@@
When $\sigma=5,n<5$, Formula (2.33), and (2.36) can be simplified as Formula (3.2), (3.2), respectively.
@@ 3.3.1 Convergent
According to Formula (3.3.1), the condition for Formula (2.30) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\sigma=1/5$, which is confirmed numerically by $\log(x|pm\delta x)$ as $P(x)\lesssim 0.20088$. Beyond the upper bound $P(x)$, the expansion is not monotonic. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.30) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.3.1), except when c is a nature number, Formula (2.33) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\sigma=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed approximately by Figure 3.3. Beyond the upper bound $P(x)$, the expansion is not monotonic. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.10) and (3.8).
@@
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}=P(x)^{2n}\sigma^{2n}\frac{2\log(2n)}{(2n)^{2}}\sigma N(\sigma);$ (3.11)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\sum_{n=1}^{\infty}P(x)^{2n}\sigma^{2n}\frac{1}{2n}\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix};$
@@ 3.3.5 Reliable
While $\delta x$ specifies the uncertainty of $x$, the calculated $\delta x$ itself can be imprecise. To guarantee $\delta x$ to be precise enough, during expansion, if the value of the variance is less than the $1/\sigma$-fold of its uncertainty, the result is invalidated as not reliable.
@@ 10.7 Source Tracing
10.8 Variable $\sigma$ Variance Arithmetic
Variance arithmetic rejects a calculation if its distributional pole or distributional zero is within the input ranges. But there are many use cases in which such intrusion cannot be avoided, so variance arithmetic of smaller $\sigma$ may be needed. It is even possible in another implementation of statistical Taylor expansion, $\sigma$ is part of a variance representation, so that a calculation closer to a distributional zero can proceed but at the expense of reducing $\sigma$. The result $\sigma$ for a calculation of imprecise inputs each with different $\sigma$ is found statistically through the corresponding $\epsilon$. Such approaches need theoretical foundation in both mathematics and statistics.

### v5

@@ 2.3 Statistical Taylor Expansion
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines bounded momentum $\zeta(n)$, in which $0<\varrho,\kappa$ are bounding factors. Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@ 2.4 Statistical Bounding
When the bounding factors $\varrho$ and $\kappa$ are limited, Formula (2.4) provides statistical bounding $x\in[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$. The probability $\epsilon$ for $x\not\in[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$ is defined as bounding leakage.
For an underlying distribution $\rho(x,\mu,\sigma)$ with known mean $\mu$ and deviation $\sigma$, when $\varrho=\kappa$, the bounding leakage $\epsilon$ is calculated as Formula (2.11) [5]. In reality, only the sample count $N$, sample mean $\overline{x}$, sample deviation $\delta x$, and the probably underlying distribution (e.g., from histogram [5]) are known after a data collection, and the bounding leakage $\epsilon$ is calculated as Formula (2.13), in which $(\mu-\overline{x})/\sigma$ is Gaussian distributed with mean 00 and deviation $\sigma/\sqrt{N}$ according to the central limit theorem [5]. When the underlying distribution is Gaussian (with the special case of $n(x)$ for normal distribution), Formula (2.11) becomes Formula (2.12), in which $\xi(\kappa)$ is the Gaussian error function [5], while Formula (2.13) becomes Formula (2.14), in which $(x-\mu)^{2}/\sigma^{2}$ is $\chi^{2}_{1}$ distributed [5], and $\sigma^{2}/(\delta x)^{2}$ is $\chi^{2}_{N-1}(\sigma)/N$ distributed (with $N-1$ degrees of freedom) [5]. Table 2.4 compares Formula (2.14) when $N=2$ with Formula (2.11). In Table 2.4, the sigma-rule $N$ is the required minimal sample count $N$ to apply the common sigma-rule of $\kappa$ for Gaussian distribution [5]. It shows that the sample count $N$ indicates the quality of the measurement.
[…]
ϵ=italic-$\epsilon=$ $\;P(\kappa<\frac{|x-\overline{x}|}{\delta x})=P(\kappa^{2}<\left(\frac{x-\overline{x}}{\sigma}\right)^{2}\frac{\sigma^{2}}{(\delta x)^{2}})$ (2.13)
[…]
$\kappa$ 1 2 3 4 5 6
[…]
Formula (2.14) $N=2$ 32.0% 15.7% 8.33% 4.55% 2.54% 1.43%
[…]
tableThe bounding leakages $\epsilon$ for different statistical bounding range $\kappa$ when the underlying uncertainty distribution is Gaussian.
@@ 2.5 Bounding Symmetry and Asymptote
A bounding $[x-\varrho\delta x,x+\kappa\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. To achieve mean-preserving, $\kappa$ determines $\varrho$. With mean-preserving bounding, Formula (2.15) and (2.16) give the result for $x\pm y$, while Formula (2.17) and (2.18) give the result for $xy$:
@@
For any input distribution $\rho(\tilde{x},x,\delta x)$ symmetric around its mean $x$, any $[x-\kappa\delta x,x+\kappa\delta x]$ bounding has property $\zeta(2n+1)=0$, which simplifies statistical Taylor expansion further.
Empirically, as Formula (2.19), (2.22), and (2.26) will demonstrate, when $n\rightarrow+\infty$, $\zeta(n)\rightarrow\kappa^{n}/n$.
@@ 2.5.1 Uniform Input Uncertainty
For uniform distribution, the bounding factor is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. According to Formula (2.19), $\zeta(2n)=(\sqrt{3})^{2n}/(2n+1)$ while $\zeta(2n+1)=0$.
@@ 2.5.2 Gaussian Input Uncertainty
If $\kappa=\infty$, $\zeta(2n)=(2n-1)!!$!, which causes Formula (2.33) for $\log(x\pm\delta x)$ and Formula (2.36) for $(x\pm\delta x)^{c}$ to diverge. When $\kappa$ has an upper bound, Formula (2.4) becomes Formula (2.20):
$\zeta(2n)=$ $2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!};$ (2.20)
$=$ $(2n-1)\zeta(\kappa,2n-2)-2N(\kappa)\kappa^{2n-1};$ (2.21)
$\kappa^{2}\ll 2n:\;\;\;\zeta(2n)\simeq 2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.22)
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$! according to Formula (2.21). When $\kappa=5,n<5$, $\left|1-\eta(2n)/(2n-1)!!\right|<10^{-3}$.
  • For large $2n$, Formula (2.20) reduces to Formula (2.22), which shows that $\zeta(2n)$ increases slower than $\kappa^{2n}$ for increasing $2n$.
@@ 2.5.3 An Input Uncertainty with Limited Range
Formula (2.23) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$. Formula (2.24) shows its mean-preserving equation. Formula (2.25) shows its bounded momentum. Formula (2.26) shows its asymptote. Formula (2.27) shows its bounding leakage.
@@
$e^{-\varrho}\varrho^{2}=e^{-\kappa}\kappa^{2};$ (2.24)
$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\kappa}\sum_{j=0}^{n+1}\frac{\kappa^{j}}{j!}\right);$ (2.25)
$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\kappa^{n+2}}{n+2}e^{-\kappa};$ (2.26)
$\epsilon$ $=1-e^{-\varrho}+e^{-\kappa};$ (2.27)
@@ 3.1 Choices of Bounding Factors
[…]
@@ 3.2 Low-Order Approximation
When $\kappa=5,n<5$, Formula (2.30), and (2.33) can be simplified as Formula (3.1), and (3.2), respectively.
@@
When $\kappa=5,n<5$, Formula (2.36), and (2.39) can be simplified as Formula (3.2), (3.2), respectively.
@@ 3.3.1 Convergent
According to Formula (3.3.1), the condition for Formula (2.33) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\kappa=1/5$, which is confirmed numerically by $\log(x|pm\delta x)$ as $P(x)\lesssim 0.20088$. Beyond the upper bound $P(x)$, the expansion is not monotonic. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.33) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.3.1), except when c is a nature number, Formula (2.36) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed approximately by Figure 3.3. Beyond the upper bound $P(x)$, the expansion is not monotonic. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.10) and (3.8).
@@
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}=P(x)^{2n}\kappa^{2n}\frac{2\log(2n)}{(2n)^{2}}\kappa N(\kappa);$ (3.11)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\sum_{n=1}^{\infty}P(x)^{2n}\kappa^{2n}\frac{1}{2n}\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix};$
@@ 3.3.5 Reliable
While $\delta x$ specifies the uncertainty of $x$, the calculated $\delta x$ itself can be imprecise. To guarantee $\delta x$ to be precise enough, during expansion, if the value of the variance is less than the $1/\kappa$-fold of its uncertainty, the result is invalidated as not reliable.
@@ 10.7 Source Tracing
10.8 Variable $\kappa$ Variance Arithmetic
As a proof of concept, the current variance arithmetic require $5\leq\kappa$ for all input data, which is impractical in reality:
  • According to Table 2.4, the required sample count for applying the sigma-rule is usually too large in reality.
  • Variance arithmetic rejects a calculation if its distributional pole or distributional zero is within the input ranges. But there are many use cases in which such intrusion cannot be avoided, so variance arithmetic of smaller $\kappa$ may be needed.
Instead of a fixed $\kappa$, a new numerical arithmetic may based on bounding leakage itself. For example, when the input uncertainty is Gaussian distributed:
  1. 1.
    Formula (2.14) calculates the bounding leakage $\epsilon$ of an input.
  2. 2.
    Formula (2.12) calculates the effective bounding range $\kappa$ from $\epsilon$.
  3. 3.
    The result of variance arithmetic is $\kappa$ dependent.
It is not clear how to calculate the result $\kappa$ or $\epsilon$ using statistical Taylor expansion, which is important for the dependency tracing when there is more than one steps from input data to final result. Such approaches need theoretical foundation in both mathematics and statistics.